OpenStudy (anonymous):

find the zeros of the functions below 1.f(x)=x^2+6x+10 2. f(x)= 3x^2-3x+9 3. f(x)= x^2-2x-9 4. f(x)= x^2+10x=5 answers only please :)

OpenStudy (anonymous):

f(x)= x ^ 2 +6 x +10=\[\rightarrow x=-6\pm2i\] 2. f (x) = 3x ^ 2-3x +9 \[\rightarrow x=(1\pm\sqrt{11} i)/2\] 3. f (x) = x ^ 2-2x-9\[\rightarrow x=1\pm\sqrt10\] 4. f (x) = x ^ 2 +10 x = 5\[\rightarrow x=-10\pm\sqrt30\]

OpenStudy (anonymous):

thankkkkk youuuu